OpenStudy (anonymous):

Simplify. (c^3)^5(d^3)^0

OpenStudy (solomonzelman):

(c^3)^5(d^3)^0 (d^3)^0 = 1 because you are taking d^3 to the zeroth power, and anything to the power of zero by definition equals 1. so (c^3)^5(d^3)^0 turns into (c^3)^5 now, lets do this, (c^3)^5= c^(3 * 5) =c^15

OpenStudy (anonymous):

Thank You!

OpenStudy (solomonzelman):

You welcome!